(a+b)^2

(a+b)^2

Grade:12

1 Answers

Nishant Vora IIT Patna
9 years ago
Hello student, Please find the answer to your question below




(a + b)2
=(a + b)(a + b)
=a2 + ab + ba + b2
=a2 + 2ab + b2

Thanks,
Regards,
Nishant Vora